OpenStudy (anonymous):

An equation is shown below: 2 over 3 multiplied by x plus 8 over 3 equals 2 to the power of x What is the solution to the equation? x = –2 x = –1 x = 1 x = 2

OpenStudy (anonymous):

@epiphi , @zoejay1

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

Can you explain ?

OpenStudy (anonymous):

23×x+83=2x does that help

OpenStudy (anonymous):

wait....

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Im sorry its x=2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

multiply everything by 3 to remove denominator 2x+23=(2x)3 12x−1+23=(2x)3 x is found on 2's so we take there exponents as the bases are the same −1+3=x x=2 you can confirm by substituting i think thats how i did it i did it in my head at first and got it wrong but this is how you get the answer

OpenStudy (anonymous):

i dont know why its doing those question marks sorry

OpenStudy (anonymous):

Okay thanks !

OpenStudy (anonymous):

Ik it always does that

OpenStudy (anonymous):

sorry if thats a little confusing

OpenStudy (anonymous):

It's not and thanks !

OpenStudy (anonymous):

alright :)

OpenStudy (anonymous):

np